Solve systems by using elimination method.
3x=3+2y
-4/3x+y=1/3

Answer by mananth(16949)

3 x + -2 y = 3 .............1
-4x/3 +y = 1/3
multiply by 3
-4x+3y=1.............2
Eliminate y
multiply (1)by 3
Multiply (2) by 2
9x-6y=9
-8x+6y=2
Add the two equations
x=11
plug value of x in (1)
3x-2y=3
33-2y=3
-2y=3-33
-2y=-30
y=15
CHECK

3 x -2 y = 3
33 -30 = 3
3 = 3